OpenStudy (anonymous):

y=x^2+x-5 y=2x-3 solve for x and y

OpenStudy (anonymous):

solve \[x^2+x-5=2x-3\]

OpenStudy (anonymous):

I get 2x=x^2+x-2 uhm then subtract x both sides and x=x^2-2 ?

OpenStudy (anonymous):

you get \[x^2-x-2=0\] then factor

OpenStudy (anonymous):

(x-2)(x+1)

OpenStudy (anonymous):

yes \[(x-2)(x+1)=0\]

OpenStudy (anonymous):

oo right has to equal zero so x values will be x=2 and x=-1?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

what do I do then?

OpenStudy (anonymous):

Im kind of lost got a final =/